OpenStudy (anonymous):

. Which is the iterative rule for the sequence: 2, 9, 16, 23, 30, …? (Points : 1)

OpenStudy (anonymous):

an = 2 + 7n?

OpenStudy (amorfide):

look at the rule you have got an=2+7n when n is 1 do you get 2? 7n is correct because the difference is 7 so when you have 7n let n=1 n=2 n=3 you get 7, 14, 21 how do you get from 7 to 2 14 to 9 21 to 16

OpenStudy (anonymous):

Ok, so an = 5 + 7n is my answer then?

OpenStudy (amorfide):

you must subtract 5 not add 5

OpenStudy (amorfide):

so your answer would be an=7n-5

OpenStudy (amorfide):

where the first term is when n=1 second term is n=2

OpenStudy (amorfide):

 since we are working out the first, second third and fourth term we will let this be n the sequence goes up in 7, so we write the seven times table how do we get from n to the number in the seven times table to get from 1 to 7 we multiply by 7 to get from 2 to 14 we multiply by 7 etc so we know we must have 7n now how do we get from 7, 14, 21, 28 to the numbers in our sequence, well we must subtract 5 since 7-5=2 14-5=9 etc so our answer is an=7n-5